If the K of a weak base is 8.7 x10, what is the pH of a 0.35 M solution of this base? Number pH= 11.75
Answer #1

B dissociates as:

B +H2O -----> BH+ + OH-

0.35 0 0

0.35-x x x

Kb = [BH+][OH-]/[B]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((8.7*10^-6)*0.35) = 1.745*10^-3

since c is much greater than x, our assumption is correct

so, x = 1.745*10^-3 M

So, [OH-] = x = 1.745*10^-3 M

use:

pOH = -log [OH-]

= -log (1.745*10^-3)

= 2.7582

use:

PH = 14 - pOH

= 14 - 2.7582

= 11.2418

Answer: 11.24